# Sample-efficient diffusion-based control of complex nonlinear systems

### v1

@@ 4 SEDC: the Proposed Method
In this section, we introduce our three key innovative designs of SEDC: Decoupled State Diffusion, Denoising Network Design of Dual Mode Decomposition, and Guided Self-finetuning.
4.1 Decoupled State Diffusion (DSD)
Decoupling Control Estimation using Inverse Dynamics. There are deep connections behind states, controls, and constraints, considering both the relationship between control and state and learning the dynamics behind state evolution. Such relation becomes more complex as the dimension of the system states goes up. However, some diffusion-based methods (e.g., DiffPhyCon (Wei et al., 2024)), which jointly diffuse over state and input, learn the relationship implicitly and may generate physically inconsistent state-control pairs that violate the underlying system dynamics. Additionally, control actions are less smooth than states, making their distribution more challenging to model (Ajay et al., 2022). Therefore, rather than jointly sampling both control signals and intermediate states using the denoising network, we choose to decouple them and diffuse only states $\mathbf{y}$, i.e.
$\mathbf{x}:=\mathbf{y}_{[0:T]}.$
Then we update the prediction of control $\mathbf{u}$ sequence by inputting the generated state trajectory to an inverse dynamic model $f_{\phi}$:
$\mathbf{u}_{t,\text{update}}^{0}=f_{\phi}(\mathbf{y}_{t}^{0},\mathbf{y}_{t+1}^{0}),$
where 00 denotes the final output of the denoising timestep from the diffusion model. We parameterize it using an Autoregressive MLP and optimize it simultaneously with the denoiser via training data. Our final optimization loss function is:
$L(\theta,\phi)$ $:=\mathbb{E}_{\mathbf{x},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f},\mathbf{\epsilon}}[||\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{x}}_{\theta}(\boldsymbol{\mathbf{x}}^{k},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f})||^{2}]$ (2)
$+\mathbb{E}_{\mathbf{y}_{t},\mathbf{u}_{t},\mathbf{y}_{t+1}}[||\mathbf{u}_{t}-f_{\phi}(\mathbf{y}_{t},\mathbf{y}_{t+1})||^{2}],$
where $\mathbf{y}_{t},\mathbf{u}_{t},\mathbf{y}_{t+1}$ are sampled from the dataset. Note that the data used to train the diffusion model can also be utilized to train $f_{\phi}$.
Cost Optimization via Gradient Guidance. After training both models, we optimize the cost function $J$ through inference-time gradient guidance. During the denoising process, we modify the sampling procedure by incorporating cost gradients:
$\mathbf{\mu}_{\theta}(\mathbf{x}^{k},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f})=$ $\frac{\sqrt{\bar{\alpha}^{k-1}}\beta^{k}}{1-\bar{\alpha}^{k}}\hat{\mathbf{x}}^{0}+\frac{\sqrt{\alpha^{k}}(1-\bar{\alpha}^{k-1})}{{1-\bar{\alpha}^{k}}}\mathbf{x}^{k}$ (3)
$-\lambda\Sigma^{k}\nabla_{\mathbf{x}^{k}}J(\hat{\mathbf{x}}^{0}(\mathbf{x}^{k})),$
where $\lambda$ controls guidance strength, $\Sigma^{k}$ is the noise scale at step $k$, $\alpha^{k}:=1-\beta^{k}\text{ and }\bar{\alpha}^{k}:=\prod_{s=1}^{k}\alpha^{s}$. Since our diffusion model operates on states only, we recover control inputs using the inverse dynamics model $f_{\phi}$ at each step. This approach enables optimization of arbitrary cost functions without model retraining, while maintaining trajectory feasibility through the learned diffusion process.
Target-conditioning as Inpainting. Modeling whether the generated trajectory accurately satisfies the initial state of $\mathbf{y}_{0}^{*}$ and the desired target state of $\mathbf{y}_{f}$ can also be regarded as a constraint satisfaction of equations, that is, the generated trajectory should contain the start and target. We adopt a more direct method to solve this: we not only input it as an additional condition to the diffusion denoising network but also treat it as an inpainting problem similar to image generation. In brief, we substitute the corresponding location in the sampled trajectories $\mathbf{x}^{k-1}\sim p_{\theta}(\mathbf{x}^{k-1}|\mathbf{x}^{k},\mathbf{y}_{0}^{*},\mathbf{y}_{f})$ with the given start and target $\mathbf{y}_{0}^{*},\mathbf{y}_{f}$ after all diffusion timesteps, analogously to observed pixels in image generation (Lugmayr et al., 2022).
4.2 Dual Mode Decomposition (DMD) for Denoising Module
In this section, we propose a design for the denoising network that decomposes the modeling of linear and nonlinear modes in the sampled trajectory by a dual-Unet architecture, as shown in Figure 1.
Our design draws inspiration from control theory. For linear systems, Yan et al. (2012) demonstrated that optimal control signals have a linear relationship with a specific linear combination $\mathbf{y}_{c}$ of initial and target states. Building upon this insight, we develop a framework where a bias-free linear layer first learns this crucial linear combination $\mathbf{y}_{c}$ from the initial state $\mathbf{y}_{0}$ and target state $\mathbf{y}_{f}$. The first UNet then learns coefficients that map $\mathbf{y}_{c}$ to control signals, establishing first-order terms, while the second UNet learns coefficients for quadratic terms. These quadratic terms, introduced through residual connections, refine the first-order approximation and enhance the network’s capacity to model complex dynamics. Note that the nonlinear terms essentially come from the nonlinearity of the dynamics.
The implementation includes several key components. The network accepts as input: (1) noisy trajectory $\mathbf{x}^{k}$ with dimension ($N$) corresponding to the network’s channel dimension; (2) initial state $\mathbf{y}_{0}$ and target state $\mathbf{y}_{f}$, which generate $\mathbf{y}_{c}$ through a bias-less linear layer; and (3) diffusion timesteps $k$, encoded via sinusoidal embedding (Ho et al., 2020) as $\mathbf{k}_{\text{emb}}$. The first UNet generates first-order coefficients to compute an initial prediction using $\mathbf{y}_{c}$. Subsequently, the second UNet combines these features to generate quadratic coefficients, producing correction terms through quadratic operations with $\mathbf{y}_{c}$. The final output $\hat{\boldsymbol{x}}^{0}$ combines these components to predict the denoised trajectory.
The architecture decomposes system dynamics into linear and nonlinear components, effectively handling complex features in nonlinear control systems while maintaining numerical stability during training. The first-order and quadratic terms based on $\mathbf{y}_{c}$ incorporate fundamental control principles into the network structure, providing effective constraints for the learning process. This structured inductive bias significantly improves the model’s data efficiency, enabling reliable control strategy learning from limited training samples.
The network performs sequential transformations on the input signals. Let $B$ denote batch size, $T$ sequence length, $C_{1}$ and $C_{2}$ feature dimensions, and $N$ the dimension of $\mathbf{y}_{c}$. The input noisy trajectory $\mathbf{x}^{k}\in\mathbb{R}^{B\times T\times N}$ and $\mathbf{y}_{c}\in\mathbb{R}^{B\times C_{1}}$ are processed through two UNets to generate first-order and quadratic predictions:
$\mathbf{C}_{1}=\text{UNet}1(\mathbf{x}^{k},\mathbf{k}{\text{emb}}),$ (4)
$\mathbf{O}_{1}=\text{reshape}(\mathbf{C}_{1})\cdot\mathbf{y}_{c},$ (5)
$\mathbf{C}_{2}=\text{UNet}_{2}([\mathbf{x}^{k},\mathbf{C}1],\mathbf{k}{\text{emb}}),$ (6)
$\mathbf{O}_{2}=\mathbf{y}_{c}^{T}\cdot\text{reshape}(\mathbf{C}_{2})\cdot\mathbf{y}_{c},$ (7)
$\hat{\mathbf{x}}^{0}=\mathbf{O}_{1}+\mathbf{O}_{2},$ (8)
where $\mathbf{C}_{1}$ produces first-order coefficients $\in\mathbb{R}^{B\times T\times(N\times C_{1})}$, $\mathbf{O}_{1}$ computes linear predictions $\in\mathbb{R}^{B\times T\times N}$, $\mathbf{C}_{2}$ generates quadratic coefficients $\in\mathbb{R}^{B\times T\times(C_{1}\times N\times C_{1})}$, and both $\mathbf{O}_{2}$ and the final output $\hat{\mathbf{x}}^{0}$ are $\in\mathbb{R}^{B\times T\times N}$. We illustrate the structural framework of the denoising network in Figure 1.
@@ 4.2 Dual Mode Decomposition (DMD) for Denoising Module
4.3 Guided Self-finetuning (GSF)
Randomly generated training data cannot guarantee coverage of optimal scenarios. To generate near-optimal controls that may deviate significantly from the training distribution. To address this limitation, we propose leveraging the model’s initially generated data (under the guidance of cost function), which naturally deviates from the training distribution toward optimality, for iterative retraining to systematically expand the exploration space. This approach maintains physical consistency by ensuring generated samples adhere to the underlying system dynamics.
Our methodology involves extracting control sequences from the generated samples (i.e., the output of inverse dynamics $\mathbf{u}_{\text{update}}^{0}$) and reintroduces it into the system to interact and generate corresponding state sequences $\mathbf{y}_{\text{update}}^{0}$. Together we add the renewed $[\mathbf{u}_{\text{update}}^{0},\mathbf{y}_{\text{update}}^{0}]$ to the retrain data pool used for a new round of fine-tuning, notably without requiring explicit system parameter identification. We iterate this process over multiple rounds specified by a hyperparameter, systematically expanding the model’s exploration space to progressively approach optimal control policy. Denote the sampling process under cost $J$’s guidance and the following interacting process as $[\mathbf{u}_{\text{update}}^{0},\mathbf{y}_{\text{update}}^{0}]=\mathcal{S}(\mathbf{x}^{K},\mathbf{y}_{0}^{*},\mathbf{y}_{f},J,\Phi)$. The process can be formulated as:
$[\mathbf{u}_{\text{update}}^{0},\mathbf{y}_{\text{update}}^{0}]=\mathcal{S}_{(\mathbf{x}^{K},\mathbf{y}_{0}^{*},\mathbf{y}_{f})\sim D}(\mathbf{x}^{K},\mathbf{y}_{0}^{*},\mathbf{y}_{f},J,\Phi),$ (9)
$D=[D,[\mathbf{u}_{\text{update}}^{0},\mathbf{y}_{\text{update}}^{0}]],$ (10)
where $D$ is the training set.
We provide the algorithm form of SEDC in Appendix 1.
@@ 5 Experiments
Experiment settings. We conducted experiments on three nonlinear systems, following the instructions in the previous works for data synthesis. These systems include: the 1-D Burgers dynamics (Hwang et al., 2022; Wei et al., 2024), which serves as a fundamental model for studying nonlinear wave propagation and turbulent fluid flow; the Kuramoto dynamics (Acebrón et al., 2005; Baggio et al., 2021; Gupta et al., 2022), which is essential for understanding synchronization phenomena in complex networks and coupled oscillator systems; and the inverted pendulum dynamics (Boubaker, 2013), which represents a classical benchmark problem in nonlinear control theory and robotic systems. For each system, we generated control/state trajectory data using the finite difference method and selected 50 trajectories as the test set. Detailed descriptions of the system dynamics equations and data synthesis procedures are provided in the appendix.
[…]
Baselines. We select the following state-of-the-art(SOTA) baseline methods for comparison. For traditional control approaches, we employ the classical PID (Proportional-Integral-Derivative) controller (Li et al., 2006), which remains widely used in industrial applications. For supervised learning ,we employ Behavioral Cloning (BC) (Pomerleau, 1988), an established imitation learning approach. In terms of reinforcement learning methods, we incorporate BPPO (Zhuang et al., 2023), a state-of-the-art algorithm. For diffusion-based methods, we include several recent prominent approaches: DecisionDiffuser (DecisionDiff) (Ajay et al., 2022), which is a SOTA classifier-free diffusion-based planner; AdaptDiffuser (Liang et al., 2023), which enhances DecisionDiffuser with a self-evolving mechanism; RDM (Zhou et al., 2024), which adaptively determines the timing of complete control sequence sampling; and DiffPhyCon (Wei et al., 2024), which is specifically designed for controlling complex physical systems. Detailed descriptions of the baselines are included in Appendix D.
@@ 5.1 Overall Control Performance
Results. In Figure 2, we compare different methods’ performance across three dynamical systems using two-dimensional coordinate plots, where proximity to the lower-left corner indicates better trade-offs between control accuracy and energy efficiency. Since unstable control can lead to system failure regardless of energy efficiency, we prioritize control accuracy and report metrics at each method’s minimum Target Loss. Our method achieves the closest position to the lower-left corner in three datasets, demonstrating best balance between accuracy and efficiency despite varying system characteristics. We achieve the best Target Loss across all systems, outperforming the best baselines by 39.5%, 49.4%, and 47.3% in Burgers, Kuramoto, and IP systems respectively. This superior performance reflects our method’s enhanced dynamics learning capability under identical conditions. For energy efficiency, our method leads in Kuramoto and IP systems while remaining competitive in Burgers, trailing AdaptDiffuser by only 1.3%.
@@
Regarding method types, traditional PID control shows the poorest performance, as system complexity exacerbates the difficulties in PID control and tuning. RL-based methods are competitive against some diffusion- but sacrifice Target Loss performance and underperform compared to Diffusion-based methods in other systems. Diffusion-based methods demonstrate superior overall performance, as they better capture long-term dependencies in system dynamics compared to traditional and RL methods, avoiding myopic failure modes and facilitating global optimization of long-term dynamics.
Due to the space limits, we present detailed numeric results in Appendix E. Moreover, we visualize the control dynamics of SEDC and SOTA baselines in Appendix G.
5.2 Sample Efficiency
Experiment settings. To evaluate the sample efficiency of diffusion-based methods, we conducted experiments on all the systems using varying proportions of the full training dataset. Specifically, we trained models using 1%, 5%, 10%, 20%, and 100% of the available data and assessed their performance using the Target Loss metric on a held-out test set.
Results. Figure 3 demonstrates our method’s superior performance in controlling Burgers and Kuramoto systems compared to state-of-the-art baselines. In all systems, our approach achieves significantly lower target loss values across all training data percentages. Most notably, with only 10% of the training data, our method attains a target loss of 1.71e-4 for Burgers, 1.12e-5 for Kuramoto, and 6.35e-4 for Inverse Pendulum, matching(-5.5% in Burgers) or exceeding(+36.4% in Kuramoto abd +1.2% in Inverse Pendulum) the performance of best baseline methods trained on the complete dataset. This indicates our method can achieve state-of-the-art performance while requiring only 10% of the training samples.
Among baselines, DiffPhyCon performed poorest due to its dual diffusion model training requirement - a challenge amplified with limited data. AdaptDiffuser surpassed DecisionDiffuser through its retraining mechanism that enhances generalization in low-data settings. Our model demonstrated superior performance through efficient dynamic learning and guided finetuning strategy.
@@ 5.2 Sample Efficiency
Table 1: Performance comparison of different ablations across multiple datasets. Target loss results with 10% and 100% training sample for each method are reported. The best, second-best and worst results of each row are highlighted in bold, underlined and italics, respectively.
System Ratio Ours Ours/DSD Ours/DMD Ours/GSF

### v2

@@ 4 SEDC: the Proposed Method
4.1 Controlling with Diffusion Models
As shown in Figure 1, SEDC reformulates the control problem as a conditional trajectory generation task. The core idea is to train a diffusion model, conditioned on the initial $\mathbf{y}_{0}^{*}$ and target $\mathbf{y}_{f}$ states, to directly generate a complete state trajectory $\mathbf{y}_{[0:T]}$. Subsequently, a corresponding control sequence $\mathbf{u}_{[0:T-1]}$ is derived from this state trajectory via a learned inverse dynamics model.
Decoupled State Diffusion (DSD). Jointly modeling the state-control distribution is highly sample-intensive and risks generating physically inconsistent trajectories. To address this, we propose Decoupled State Diffusion (DSD). As shown in Figure 1, we confine the diffusion process to the state trajectory $\mathbf{y}$ alone (i.e., $\mathbf{x}:=\mathbf{y}_{[0:T]}$), as state evolution is generally smoother and more structured. The corresponding control sequence is then derived from the generated state transitions via a separately trained inverse dynamics model, $f_{\phi}$: $\tilde{\mathbf{u}}_{t,\text{update}}^{0}=f_{\phi}(\mathbf{y}_{t}^{0},\mathbf{y}_{t+1}^{0}),$ where the superscript 0 denotes the final denoised output of the diffusion model.
A common concern with deterministic inverse dynamics models $f_{\phi}$ is the potential for averaging valid controls in multi-modal scenarios (where distinct actions yield identical transitions), leading to physically invalid actions. However, Unlike joint modeling state and control, which risks generating out-of-distribution pairs, DSD ensures that the input to the inverse model $(\mathbf{y}_{t},\mathbf{y}_{t+1})$ is always drawn from a valid, coherent manifold generated by the diffusion model (Chen et al., 2025). Even in settings with control non-uniqueness, the inverse model is only tasked with finding a feasible control that satisfies this valid transition, a significantly well-posed problem compared to unconstrained joint generation. We also evaluate and discuss SEDC’s robustness to non-invertibility in Section 6.
We then optimize $f_{\phi}$ simultaneously with the denoiser. The loss function is:
[…]
where $\mathbf{y}_{t},\mathbf{u}_{t},\mathbf{y}_{t+1}$ are sampled from the dataset. After training, the denoising process generates a state trajectory $\mathbf{y}_{0:T}$ that is both physically plausible and conditioned on the start and target states. To further refine this process, we introduce two key mechanisms: inpainting for hard constraint satisfaction and gradient guidance for soft optimization.
Target-Conditioning via Inpainting. To ensure the generated trajectory strictly adheres to the given initial state $\mathbf{y}_{0}^{*}$ and target state $\mathbf{y}_{f}$, we treat the problem as a form of trajectory inpainting. While these states are provided as conditions to the denoising network, we enforce them as hard constraints during the sampling process. Specifically, at each denoising step $k$, after sampling a potential trajectory $\mathbf{x}^{k-1}\sim p_{\theta}(\mathbf{x}^{k-1}|\mathbf{x}^{k},\mathbf{y}_{0}^{*},\mathbf{y}_{f})$, we replace its start and end points with the ground truth values: $\mathbf{x}^{k-1}_{0}\leftarrow\mathbf{y}_{0}^{*}$ and $\mathbf{x}^{k-1}_{T}\leftarrow\mathbf{y}_{f}$. This technique, analogous to inpainting in image generation (Lugmayr et al., 2022), guarantees that the final output satisfies the boundary conditions.
Cost Optimization via Gradient Guidance. Beyond satisfying the boundary constraints, we aim to find a trajectory that minimizes a given cost function $J$ (e.g., control energy). We achieve this through inference-time gradient guidance. Building upon the inpainting-enforced sampling, we further modify the mean of the denoising distribution by incorporating the cost gradient:
$\mathbf{\mu}_{\theta}(\mathbf{x}^{k},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f})=$ $\frac{\sqrt{\bar{\alpha}^{k-1}}\beta^{k}}{1-\bar{\alpha}^{k}}\hat{\mathbf{x}}^{k}+\frac{\sqrt{\alpha^{k}}(1-\bar{\alpha}^{k-1})}{{1-\bar{\alpha}^{k}}}\mathbf{x}^{k}$ (3)
$-\lambda\Sigma^{k}\nabla_{\mathbf{x}^{k}}J(\hat{\mathbf{x}}^{k}(\mathbf{x}^{k})),$
where $\lambda$ controls guidance strength and the superscript $k$ of $\hat{\mathbf{x}}^{k}:=\hat{\mathbf{y}}_{[0:T]}^{k}$ denotes the clean output from the denoiser at step $k$. Since our diffusion model operates on states only, we recover the control sequence $\tilde{\mathbf{u}}_{t}^{k}=f_{\phi}(\hat{\mathbf{y}}_{t}^{k},\hat{\mathbf{y}}_{t+1}^{k})$ to compute the cost $J({\mathbf{u}})$ at each step. This combined approach of inpainting and guidance ensures the final generated trajectory is not only feasible and satisfies hard constraints but is also optimized for the desired cost objective.
4.2 Dual-Mode Decomposition (DMD) for Denoiser
[…]
This decomposition imposes a strong inductive bias: modeling a dominant linear part and a subtle nonlinear correction is a more stable and sample-efficient task than forcing a monolithic network to learn the entire complex function from scratch. We implement this with a dual-UNet architecture, as illustrated in Figure 1. The network is conditioned on a vector $\mathbf{y}_{c}=f_{c}([\mathbf{y}_{0},\mathbf{y}_{f}])\in\mathbb{R}^{B\times C_{1}}$, which is generated from the start and target states via a bias-less linear layer $f_{c}$.
Given the noisy trajectory $\mathbf{x}^{k}\in\mathbb{R}^{B\times T\times N}$ and time embedding $\mathbf{k}_{\text{emb}}$, the first stage produces a linear prediction $\mathbf{O}_{1}\in\mathbb{R}^{B\times T\times N}$ from first-order coefficients $\mathbf{C}_{1}$:
$\mathbf{C}_{1}=\text{UNet}_{1}(\mathbf{x}^{k},\mathbf{k}_{\text{emb}}),\quad\mathbf{O}_{1}=\text{reshape}(\mathbf{C}_{1})\cdot\mathbf{y}_{c}.$ (4)
In the second stage, UNet2 combines the input $\mathbf{x}^{k}$ and intermediate features $\mathbf{C}_{1}$ to generate quadratic coefficients $\mathbf{C}_{2}$, which yield a nonlinear correction term $\mathbf{O}_{2}\in\mathbb{R}^{B\times T\times N}$:
$\mathbf{C}_{2}=\text{UNet}_{2}([\mathbf{x}^{k},\mathbf{C}_{1}],\mathbf{k}_{\text{emb}}),\quad\mathbf{O}_{2}=\mathbf{y}_{c}^{T}\cdot\text{reshape}(\mathbf{C}_{2})\cdot\mathbf{y}_{c}.$ (5)
The final denoised prediction is the sum of these components:
$\hat{\mathbf{x}}^{0}=\mathbf{O}_{1}+\mathbf{O}_{2}\in\mathbb{R}^{B\times T\times N}.$ (6)
Here, $\mathbf{C}_{1}\in\mathbb{R}^{B\times T\times(N\times C_{1})}$ and $\mathbf{C}_{2}\in\mathbb{R}^{B\times T\times(C_{1}\times N\times C_{1})}$ represent the learned coefficient tensors.
4.3 Guided Self-finetuning (GSF)
Randomly generated training data cannot guarantee coverage of optimal scenarios. To generate near-optimal controls that may deviate significantly from the training distribution, we propose leveraging the model’s initially generated data (under the guidance of the cost function), which naturally deviates from the training distribution toward optimality, for iterative retraining to systematically expand the manifold coverage. This approach maintains physical consistency by ensuring generated samples adhere to the underlying system dynamics.
As shown in Figure 1, our methodology involves extracting control sequences from the generated samples (i.e., the output of inverse dynamics $\tilde{\mathbf{u}}^{0}$) and reintroducing them into the ground-truth simulator (or the physical environment if accessible) to generate corresponding state sequences $\tilde{\mathbf{y}}^{0}$. While this introduces an online interaction component, it is highly practical in scientific control tasks (e.g., fluid dynamics) where forward simulation is possible but computationally expensive to run exhaustively. GSF serves as a sample-efficient bridge between offline initialization and optimal control. Together, we add the renewed $\tilde{\mathbf{u}}^{0},\tilde{\mathbf{y}}^{0}]$ to the retrain data pool used for a new round of fine-tuning, notably without requiring explicit system parameter identification. We iterate this process over multiple rounds specified by a hyperparameter, systematically expanding the model’s manifold support to progressively approach the optimal control law. Denote the sampling process under cost $J$’s guidance and the following interacting process as $\tilde{\mathbf{u}}^{0},\tilde{\mathbf{y}}^{0}]=\mathcal{S}(\mathbf{x}^{K},\mathbf{y}_{0}^{*},\mathbf{y}_{f},J,\Phi)$. The process can be formulated as:
$\tilde{\mathbf{u}}^{0},\tilde{\mathbf{y}}^{0}]=\mathcal{S}_{(\mathbf{x}^{K},\mathbf{y}_{0}^{*},\mathbf{y}_{f})\sim D}(\mathbf{x}^{K},\mathbf{y}_{0}^{*},\mathbf{y}_{f},J,\Phi),$ (7)
$D=[D,[\tilde{\mathbf{u}}^{0},\tilde{\mathbf{y}}^{0}]],$ (8)
where $D$ is the training set and $\mathcal{S}(\mathbf{x}^{K},\mathbf{y}_{0}^{*},\mathbf{y}_{f},J,\Phi)$ denotes the full process of generating a guided trajectory using cost $J$ and then interacting with the system dynamics $\Phi$ to get the corresponding state sequence. We provide the algorithm form of SEDC in Appendix A.
@@ 5 Experiments
5.1 Experiment Settings
We conducted experiments on multiple nonlinear physics systems, following the instructions in the previous works for data synthesis. These systems include: the 1-D Burgers dynamics (Hwang et al., 2022; Wei et al., 2024b), which is a high-dimensionality (128-dim, significantly larger than standard low-dimensional robotics benchmarks) dynamic system for studying nonlinear wave propagation and turbulent fluid flow; the Kuramoto dynamics (Acebrón et al., 2005; Baggio et al., 2021; Gupta et al., 2022), which is essential for understanding synchronization phenomena in complex networks and coupled oscillator systems; the inverted pendulum dynamics (Boubaker, 2013), a classic nonlinear control baseline. To further assess the scalability of our approach in extreme high-dimensional settings, we also evaluate SEDC on Jellyfish locomotion, a challenging 2D PDE control task involving complex fluid-solid interactions (detailed in Appendix G.2). For each system, we generated control/state trajectory data using the finite difference method and selected 50 trajectories as the test set. We assume full state observability throughout the experiments. Detailed descriptions of the system dynamics equations and data synthesis procedures are provided in Appendix B. Implementation details and time consumption reports are provided in Appendix C.
We evaluate two metrics which is crucial in complex system control: Target Loss, the mean-squared-error (MSE) of $\mathbf{y}_{T}$ and desired target $\mathbf{y}_{f}$, i.e. $\frac{1}{N}\|\mathbf{y}_{T}-\mathbf{y}_{f}\|^{2}$. (Note that $\mathbf{y}_{T}$ is obtained by simulating the real system using the control inputs generated by each method, along with the given initial state conditions, rather than extracted from the sample trajectories of the diffusion-based methods); Energy $J=\int_{0}^{T}|\mathbf{u}(t^{\prime})|^{2}dt^{\prime}$, which measures the cumulative control effort required to achieve the target state. Lower values of both metrics indicate better performance.
Baselines. We evaluate SEDC against two categories of control methods. Iterative Methods: we include the classical PID controller (Li et al., 2006) and data-driven sequential learning approaches, including Behavioral Cloning (BC) (Pomerleau, 1988) and Batch Proximal Policy Optimization (BPPO) (Zhuang et al., 2023). Global Methods: we also compare against state-of-the-art diffusion-based methods that generate holistic plans: DecisionDiffuser (DecisionDiff) (Ajay et al., 2022), AdaptDiffuser (AdaptDiff) (Liang et al., 2023), RDM (Zhou et al., 2024), and DiffPhyCon (Wei et al., 2024b). These global methods avoid error accumulation by solving for the entire trajectory holistically. Detailed descriptions are included in Appendix D.
@@ 5.2 Overall Control Performance
In Figure 2, we compare different methods’ performance across three dynamical systems using two-dimensional coordinate plots, where proximity to the lower-left corner indicates better trade-offs between control accuracy and energy efficiency. Since unstable control can lead to system failure regardless of energy efficiency, we prioritize control accuracy and report metrics at each method’s minimum Target Loss.
SEDC consistently achieves state-of-the-art performance. Our method secures the position closest to the origin in all three datasets, demonstrating the best balance between accuracy and efficiency. Specifically, SEDC achieves the lowest Target Loss across all systems, outperforming the strongest baselines by 39.5%, 49.4%, and 47.3% in the Burgers, Kuramoto, and IP systems, respectively. This highlights its superior capability in learning complex dynamics. In terms of energy cost, SEDC leads on the Kuramoto and IP systems and remains highly competitive on the Burgers system.
@@
Analysis of Baselines. The results reveal clear performance tiers among different method families. Traditional PID control shows the poorest performance, as system complexity exacerbates the difficulties in PID control and tuning. Sequential learning methods (e.g., BPPO) are competitive against some diffusion-based approaches but sacrifice Target Loss performance and underperform compared to global trajectory planning methods in other systems. Diffusion-based methods demonstrate superior overall performance, as they better capture long-term dependencies in system dynamics compared to traditional and iterative sequential methods, avoiding myopic failure modes and facilitating global optimization of long-term dynamics.
Due to space constraints, Appendices E and H provide detailed numerical results, including standard errors, Pareto frontier comparisons, and control dynamics visualizations. A comparison with a data-driven MPC in Appendix G.1 further underscores the advantages of our global planning approach. While our main experiments assume full-state observability, we validate SEDC’s robustness to significant observation noise in Appendices G.6. Furthermore, Appendices G.2 validate the method’s effectiveness on higher-dimensional physics systems.
@@
Table 1: Ablation performance across datasets. Target Loss at 10% and 100% training data is reported. Best/second-best/worst in each row are in bold/underlined/italics.
System Ratio Ours w/o DSD w/o DMD w/o GSF
@@
5.3 Sample Efficiency
To evaluate the sample efficiency of diffusion-based methods, we conducted experiments on all the systems using varying proportions of the full training dataset. Specifically, we trained models using 1%, 5%, 10%, 20%, and 100% of the available data and assessed their performance using the Target Loss metric on a held-out test set. Figure 3 demonstrates our method’s superior performance in controlling Burgers and Kuramoto systems compared to state-of-the-art baselines. In all systems, our approach achieves significantly lower target loss values across all training data percentages. Most notably, with only 10% of the training data, our method attains a target loss of 1.71e-4 for Burgers, 1.12e-5 for Kuramoto, and 6.35e-4 for Inverse Pendulum, matching(-5.5% in Burgers) or exceeding(+36.4% in Kuramoto and +1.2% in Inverse Pendulum) the performance of best baseline methods trained on the complete dataset. This indicates our method can achieve state-of-the-art performance while requiring only 10% of the training samples.
Among the baselines, DiffPhyCon’s complex training requirement (Sec. 2) makes it particularly sample-inefficient. While AdaptDiffuser’s self-tuning improves upon DecisionDiffuser, SEDC’s advantage stems from its efficient dynamics learning and a filter-free finetuning strategy. Unlike AdaptDiffuser, our GSF mechanism integrates all guided trajectories without discriminator filtering, promoting data diversity and a better exploration-exploitation balance.
(a) .
(b) .
Figure 4: Ablation analysis of SEDC components. (a) Target Loss decrease rate of Decoupled State Diffusion (DSD), measured as the reduction relative to w/o DSD across state dimensions in Burgers and Kuramoto. (b) Energy cost ($J$) reduction over Guided Self-finetuning (GSF) rounds, with most gains in the first round and subsequent rounds indicating convergence.